A company's warehouse contents were destroyed by a flood on September 12. The following information was the only information that was salvaged:

  1. Inventory, beginning: $29,000
  2. Purchases for the period: $18,000
  3. Sales for the period: $56,000
  4. Sales returns for the period: $800

   
The company's average gross profit ratio is 30%. What is the estimated cost of the lost inventory?
Answer #1

Cost of Goods available for sale = Beg inventory + Purchases

= 29,000+18,000

= 47,000

Cost of Goods Sold = Net sales - Gross profit

= (56,000-800) - 30%

= 38,640.

Estimated cost of lost inventory = 47,000 - 38,640

= 8360